OpenStudy (anonymous):

Darlene Ramirez bought a home for $140,000. She put 20% down with a mortgage rate of 7.5% for 25 years. Her yearly payments are:

OpenStudy (amistre64):

so how much does she have to borrow if shes paid 20% upfront?

OpenStudy (amistre64):

also its an odd statement, early payments. most mortagages are a monthly payment, so would they want total payments made per year? or are we determining that a payment is made once a year?

OpenStudy (amistre64):

**yearly payments is an odd statement ....

OpenStudy (anonymous):

it is asking for the payment made once a year

OpenStudy (amistre64):

ok ... thats doable

OpenStudy (amistre64):

do you have a formula you use? or should we make one that I made up myself?

OpenStudy (amistre64):

given a number of payment periods, n, for a given loan amount B, and a stated periodic compounding rate, k, the balance of the loan at any point in time is determined as: \[B_n=Bk^n-P\frac{k^n-1}{k-1}\] the loan is paid when Bn = 0, so we can solve for P, payments \[0=Bk^n-P\frac{k^n-1}{k-1}\] \[Bk^n=P\frac{k^n-1}{k-1}\] \[P=Bk^n\frac{k-1}{k^n-1}\] which we can reduce to the textbook formula but i find this version more useful

OpenStudy (amistre64):

\[P=140000(1+.075)^{25}\frac{(1+.075)-1}{(1+.075)^{25}-1}\]

OpenStudy (amistre64):

forgot to remove the 20% down payent of course but thats just an adjustment to B

OpenStudy (amistre64):

we are borrowing 80% of 140 000 so multiply the results by .80

OpenStudy (anonymous):

so we don't use the 20% down payment?

OpenStudy (anonymous):

so we don't use the 20% down payment?

OpenStudy (amistre64):

we did we paid in 20% of 140000 so the amount needed to borrow is the remaining 80% of 140000 our original balance is therefore just 140000(.80)

OpenStudy (amistre64):

we can either figure that number out, or just adjust the balance as is by multiplying it by .80

OpenStudy (amistre64):

\[P=140000(\color{red}{.80})(1+.075)^{25}\frac{(1+.075)-1}{(1+.075)^{25}-1}\]